OpenStudy (anonymous):

My question is to find the derivative of y=(11/3)cos^-1(5-2x) I have y'=(11.3)(-1/sqrt(1-(5-2x^2))*2 I don't know what to do next please help

OpenStudy (anonymous):

My question is to find the derivative of y=(11/3)cos^-1(5-2x) I have y'=(11/3)(-1/sqrt(1-(5-2x^2))*2 I don't know what to do next please help

OpenStudy (anonymous):

From what I see the general form of your derivative is correct except that last 2 I believe should be a negative. I would also guess that the (5-2x) is squared in your notation, you just mistyped. You can simplify it if you want, but I'd say that's fine. Depends on the teacher.